OpenStudy (anonymous):

Solve the equation. Show your work. 8x – 12 = 4x + 24

OpenStudy (anonymous):

8x-4x=24+12 4x=36 x=9

OpenStudy (nubeer):

subtract 4x on both sides 8x - 4x -12 = 4x -4x +24 4x -12 = 24 4x =24-12.. i believe u can do frm here..

OpenStudy (anonymous):

U should move the 8x to the right side and =24 o the left side then u will be able to solve it... :)

OpenStudy (anonymous):

show me

OpenStudy (anonymous):

-12-24=4x-8x -36=-4x x=9 Same result, just more -....

OpenStudy (anonymous):